A child whose weight is 270 N slidesdown a 6.2 m playground slide that makesan angle of 20° with the horizontal. The coefficient of kineticfriction is 0.10.
(a) How much energy is transferred to thethermal energy?
1 J

(b) If she starts at the top with a speed of 4.97 m/s, what is herspeed at the bottom?
2 m/s
Answer #1
Consider



W=mg= 270N
f= μmgcos(20)
(a) How much energy is transferred to the thermal energy?
W thermal = fL= μmgcos(20) L
W thermal = fL= 0.10 x 270 cos(20) x 6.3
W thermal = 160 J

(b) If she starts at the top with a speed of 4.97 m/s, what is herspeed at the bottom?

Pe1 + Ke1-Wf=Kef  ( the loss due to friction was not figuredin)
mgh + 0.5mV12 -μmgcos(20) L =0.5mVf2

Vf2 = 2gh +  V12 - μgcos(20) L
Vf = √(2gh +   V12- μgcos(20) L) whereh = L sin(20)
Vf = √(2g L sin(20) +  V12    -μgcos(20) L)
Vf = √(2x 9.81 x 6.3 sin(20)+   4.972- 0.1 x9.81 cos(20) 6.3)
Vf = 7.8m/s
Answer #2

ANSWER :


1. 


(a)


Normal reaction = mg cos(20) = 270 cos(20)

Frictional force = µ * 270 cos(20) = 0.10 * 270 cos(20)

W. D against friction = 27 cos(20) * 6.2   (J)


Energy transferred to thermal energy 

= WD against friction 

= 27 cos(20) * 6.2

= 157.305 J (ANSWER).


(b)


K. E at the top = 1/2 m v^2 = 1/2 W/g v^2 = 1/2 * 270/9.8 * (4.7)^2 = 304.30 J

Height of slide top = 6.2 sin(20) = 2.12 m

P. E at the top of slide = W * h = 270 * 2.12 = 572.4 J


K. E. At the bottom 

= K. E at the top + P.E. at the top - Thermal energy generated

= 304.30 + 572.40 - 157.305

= 719.395 J


Let her speed at the bottom be V m/sec.

So,


1/2 W/g V^2 = 719.395

=> V^2 = 719.395* 2g / W = 719.397 * 2*9.8 / 270 = 52.223

=> V = sqrt(52.223)

=> V = 7.227 m/sec 

Her speed at the bottom of slide is = 7.223 m/sec (ANSWER).


answered by: Tulsiram Garg

> In the last line. correct , " 7.223 " as " 7.227 "

Tulsiram Garg Sun, Oct 10, 2021 11:17 AM